Star Company paid $5 per share dividend a week ago. General expectation is that the dividend would grow at 8% per year indefinitely.
a) What is the expected dividend next year?
b) If an investor requires 12%, what would be the price of the share?
1
Expert's answer
2015-09-09T00:00:43-0400
Star Company paid $5 per share dividend a week ago. General expectation is that the dividend would grow at 8% per year indefinitely.
a) The expected dividend next year will be 5*1.08 = $5.
b) If an investor requires 12%, the price of the share will be P = D/(r - i) = 5/(0.12 - 0.08) = $125.